OpenStudy (chloe103):

The masses of four planets in a solar system outside the Milky Way are given below. Mass of Planet A: 5.97 x 10^24 kg; Mass of Planet B: 3.30 x 10^23 kg; Mass of Planet C: 1.89 x 10^27 kg; Mass of Planet D: 4.87 x 10^24 kg What is the order of these planets from least to greatest mass? Planet C; Planet;B Planet D; Planet A Planet C; Planet A;Planet D; Planet B Planet B; Planet D;Planet A; Planet C Planet A;Planet D; Planet B; Planet C

OpenStudy (triciaal):

one approach use the same base unit then find the equivalent value put the values in numerical order from least to greatest read the correct option that matches your result

OpenStudy (triciaal):

I would use 10^23 as the base unit A = 59.7 x10^23 B = 3.3 x10^23 C = greatest D = 48.7 x 10^23 so 3.3, 48.7, 59.7 then C matching option C

OpenStudy (iiomqkinqx3):

the answer is panet D JUST HOW BIG MY D IS flutterCKCKC